What is the area of the sector, in square units, determined by an arc with a measure 45° in a circle with radius 6? Round to the nearest tenth.

help please thanks.! :)

Answer by lwsshak3(11628)

What is the area of the sector, in square units, determined by an arc with a measure 45° in a circle with radius 6? Round to the nearest tenth
***
Area of a circle=πr^2
Area of sector=45/360*π6^2=14.14 sq units